# New Borwein-type conjectures

### v1

@@ Abstract.
Motivated by recent research of Krattenthaler and Wang, we propose five new “Borwein-type” conjectures modulo $3$ and two new “Borwein-type” conejctures modulo $5$.
@@
Borwein conjectures, Borwein-type conjectures, sign-pattern.
@@ 1. Introduction
In $1990$, Peter Borwein unearthed several mysteries involving sign patterns of coefficients of certain polynomials. He then personally communicated George Andrews who stated them as the following in his paper [1]:
Conjecture 1.1.
(First Borwein Conjecture) For the polynomials $A_{n}(q)$, $B_{n}(q)$, and $C_{n}(q)$ defined by
@@
Conjecture 1.2.
(Second Borwein Conjecture) For the polynomials $\alpha_{n}(q)$, $\beta_{n}(q)$, and $\gamma_{n}(q)$ defined by
@@
Conjecture 1.3.
(Third Borwein Conjecture) For the polynomials $\nu_{n}(q)$, $\phi_{n}(q)$, $\chi_{n}(q)$, $\psi_{n}(q)$, and $\omega_{n}(q)$ defined by
@@
Although Andrews did not prove any of the above conjectures, he stated [1, eq. $(1.2)$] a partition theoretic analogue of Conjecture 1.1 and also derived explicit sum representations of the polynomials $A_{n}(q)$, $B_{n}(q)$, and $C_{n}(q)$. It was recently proved analytically by Chen Wang using saddle point method [2]. Apart from other tools, his proof relied on a theorem of Andrews [1, Theorem $4.1$] and the following three recursive relations also given by Andrews [1, eq. $(3.1)$ $(3.3)$]:
@@
Recently, Christian Krattenthaler and Chen Wang [3] proved Conjecture 1.1 and Conjecture 1.2 completely. They also gave “two thirds” proof of the following cubic version which was missed by both Borwein and Andrews and can be stated as follows:
Conjecture 1.4.
(Cubic Borwein Conjecture) For the polynomials $\kappa_{n}(q)$, $\delta_{n}(q)$, and $\theta_{n}(q)$ defined by
@@
Unlike Wang [2] who considered Andrews’ sum representations for his proof, Krattenthaler and Wang [3] applied the saddle point method directly to the “First Borwein polynomial” $\prod\limits_{j=1}^{n}(1-q^{3j-2})(1-q^{3j-1})$ and its higher powers. They also provided ideas [3, Conjecture $11.1$] for (at least) “three fifths” proof of Conjecture 1.3 when the modulus $3$ is replaced by $5$.
@@
For positive integers $k$ and $i$, we consider polynomials of the form
$P_{n,i}^{(k)}(q):=\dfrac{(q;q)_{kn}^{i}}{(q^{k};q^{k})_{n}^{i}}.$
For $k=7$ above, Krattenthaler and Wang [3, Conjecture $11.5$] stated an interesting modulus $7$ Borwein Conjecture which is as follows:
Conjecture 1.5.
(Modulus $7$ Borwein Conjecture) For positive integers $n$, consider the expansion of the polynomial
$P_{n,1}^{(7)}(q):=\dfrac{(q;q)_{7n}}{(q^{7};q^{7})_{n}}=\sum\limits_{m=0}^{21n^{2}}d_{m}(n)q^{m}.$
@@
Motivated by Conjecture 1.5, in Section 2, we state new “Borwein-type” sign-pattern conjectures for the polynomial $P_{n,i}^{(k)}(q)$ where $(k,i)\in\{(3,4),(3,5),(3,6),(3,7),(3,8),(5,2),(5,3)\}$. We conclude with some possible future directions in Section 3.
@@ 2. New Conjectures
Let $n$ be a positive integer and $i\in\{4,5,6,7,8\}$. Consider the expansion of the polynomial $P_{n,i}^{(3)}(q)$ defined by
$P_{n,i}^{(3)}(q):=\dfrac{(q;q)_{3n}^{i}}{(q^{3};q^{3})_{n}^{i}}=\sum\limits_{m=0}^{3in^{2}}c_{m}^{(i)}(n)q^{m}.$
@@
where $\alpha^{(i)}(n)$ seems to stabilise around according to the following table
value of $i$ limiting value of $\alpha^{(i)}(n)$
$4$ $0.751$
$5$ $1.270$
$6$ $1.778$
$7$ $2.282$
$8$ $2.784$
@@
Since the polynomial $P_{n,i}^{(3)}(q)$ is palindromic for any positive integer $i$, we can also predict the signs of the coefficients $c_{3m+1}^{(i)}(n)$ in Conjecture 2.1. This is true because of the following duality result which is easy to prove.
Theorem 2.2.
Consider the polynomial $P_{n,i}^{(3)}(q)$ as defined in Conjecture 2.1. If we write
$P_{n,i}^{(3)}(q)=X_{n,i}(q^{3})+qY_{n,i}(q^{3})+q^{2}Z_{n,i}(q^{3}),$
then the following relations hold
$X_{n,i}(q)=q^{in^{2}}X_{n,i}(1/q)$
@@
Let $n$ be a positive integer and $i\in\{2,3\}$. Consider the expansion of the polynomial $P_{n,i}^{(5)}(q)$ defined by
$P_{n,i}^{(5)}(q):=\dfrac{(q;q)_{5n}^{i}}{(q^{5};q^{5})_{n}^{i}}=\sum\limits_{m=0}^{10in^{2}}d_{m}^{(i)}(n)q^{m}.$
@@
where $\tilde{\alpha}^{(i)}(n)$ and $\tilde{\beta}^{(i)}(n)$ seem to stabilise around according to the following table
value of $i$ limiting value of $\tilde{\alpha}^{(i)}(n)$ limiting value of $\tilde{\beta}^{(i)}(n)$
$2$ $1.133$ $1.011$
$3$ $2.132$ $2.001$
@@
Since the polynomial $P_{n,i}^{(5)}(q)$ is palindromic for any positive integer $i$, we can also predict the signs of the coefficients $d_{5m+1}^{(i)}(n)$ and $d_{5m+2}^{(i)}(n)$ in Conjecture 2.3. This is true because of the following duality result which we leave as an exercise for the interested reader to prove.
Theorem 2.4.
Consider the polynomial $P_{n,i}^{(5)}(q)$ as defined in Conjecture 2.3. If we write
$P_{n,i}^{(5)}(q)=A_{n,i}(q^{5})+qB_{n,i}(q^{5})+q^{2}C_{n,i}(q^{5})+q^{3}D_{n,i}(q^{5})+q^{4}E_{n,i}(q^{5}),$
then the following relations hold
$A_{n,i}(q)=q^{2in^{2}}A_{n,i}(1/q),$
$D_{n,i}(q)=q^{2in^{2}-1}C_{n,i}(1/q),$
@@
It must be noted that Krattenthaler and Wang [3, Remark $11.6.(2)$] mention that for $\delta\in\{2,3\}$, the coefficient of $q^{5m}$ in $(q;q)_{5n}^{\delta}/(q^{5};q^{5})_{n}^{\delta}$ is non-negative for all $m$, however, they don’t find the limiting values of $\tilde{\alpha}^{(\delta)}(n)$ and $\tilde{\beta}^{(\delta)}(n)$.
@@ 3. Concluding Remarks
    It would be very interesting to investigate whether the asymptotic approach of Krattenthaler and Wang in [3] applies to Conjecture 2.1 and Conjecture 2.3.
@@
    The sign-pattern of coefficients is much more complicated for $P_{n,i}^{(3)}(q)$ when $i\geq 9$ and for $P_{n,i}^{(5)}(q)$ when $i\geq 4$ but is definitely something worth exploring more.

### v3

@@ Abstract.
Motivated by recent research of Wang and Krattenthaler, we use Maple to propose five new “Borwein-type” conjectures modulo $3$ and two new “Borwein-type” conjectures modulo $5$.
@@
Borwein conjectures, Borwein-type conjectures, sign-pattern, $q$-series.
@@ 1. Introduction
In $1990$, Peter Borwein unearthed several mysteries involving sign patterns of coefficients of certain polynomials. He then personally communicated his observations to George Andrews who stated them as the following in his paper [1]:
Conjecture 1.1 (First Borwein Conjecture).
For the polynomials $A_{n}(q)$, $B_{n}(q)$, and $C_{n}(q)$ defined by
@@
Conjecture 1.2 (Second Borwein Conjecture).
For the polynomials $\alpha_{n}(q)$, $\beta_{n}(q)$, and $\gamma_{n}(q)$ defined by
@@
Conjecture 1.3 (Third Borwein Conjecture).
For the polynomials $\nu_{n}(q)$, $\phi_{n}(q)$, $\chi_{n}(q)$, $\psi_{n}(q)$, and $\omega_{n}(q)$ defined by
@@
Although Andrews did not prove any of the above conjectures, he derived explicit sum representations [1, Theorem $3.1$, eq. $(3.4)$$(3.6)$] of the polynomials $A_{n}(q)$, $B_{n}(q)$, and $C_{n}(q)$ respectively. Conjecture 1.1 was recently proved analytically by Chen Wang using the saddle point method [2]. Apart from other tools, his proof relied on a theorem of Andrews [1, Theorem $4.1$] and the following three recursive relations also given by Andrews [1, eq. $(3.1)$$(3.3)$]:
@@
Recently, Chen Wang and Christian Krattenthaler [3] proved Conjecture 1.1 and Conjecture 1.2. Unlike Wang [2] who considered Andrews’ sum representations for his proof, Wang and Krattenthaler [3] applied the saddle point method directly to the “First Borwein polynomial” $\prod\limits_{j=1}^{n}(1-q^{3j-2})(1-q^{3j-1})$ and its higher powers. They also gave [3, Conjecture $1.3$] a cubic conjecture which was missed by both Borwein and Andrews and can be stated as follows:
Conjecture 1.4 (Cubic Borwein Conjecture).
For the polynomials $\kappa_{n}(q)$, $\delta_{n}(q)$, and $\theta_{n}(q)$ defined by
@@
They proved “two thirds” of Conjecture 1.4 above by showing that all coefficients of $\kappa_{n}(q)$ are non-negative, “one half” of the coefficients of $\delta_{n}(q)$ are non-negative, and “one half” of the coefficients of $\theta_{n}(q)$ are non-negative. They also provided ideas [3, Conjecture $11.1$] along similar lines for (at least) “three fifths” of a proof of Conjecture 1.3 when the modulus $3$ is replaced by $5$.
@@
For positive integers $n$ and $k$, we consider polynomials of the form
$P_{n,k}(q):=\dfrac{(q;q)_{kn}}{(q^{k};q^{k})_{n}}=\prod\limits_{j=1}^{k-1}(q^{j};q^{k})_{n}.$
For $k=7$ above, Wang and Krattenthaler [3, Conjecture $11.5$] stated an interesting modulus $7$ Borwein Conjecture which is as follows:
Conjecture 1.5 (Modulus $7$ Borwein Conjecture).
For positive integers $n$, consider the expansion of the polynomial
$P_{n,7}(q):=\dfrac{(q;q)_{7n}}{(q^{7};q^{7})_{n}}=\sum\limits_{m=0}^{21n^{2}}d_{m}(n)q^{m}.$
@@
Motivated by Conjecture 1.5, in Section 2, we state new “Borwein-type” sign-pattern conjectures for the polynomial $P_{n,k}^{\,i}(q)$ where $(k,i)\in\{(3,4),(3,5),(3,6),(3,7),(3,8),(5,2),(5,3)\}$. We conclude with some possible future directions in Section 3.
@@ 2. New Conjectures
Let $n$ be a positive integer and $i\in\{4,5,6,7,8\}$. Consider the expansion of the polynomial $P_{n,3}^{\,i}(q)$ defined by
$P_{n,3}^{\,i}(q):=\dfrac{(q;q)_{3n}^{i}}{(q^{3};q^{3})_{n}^{i}}=\sum\limits_{m=0}^{3in^{2}}c_{m}^{(i)}(n)q^{m}.$
@@
where $\alpha^{(i)}(n)$ tends to a limiting value as $n$ tends to $\infty$ according to the following table
value of $i$ $\lim\limits_{n\rightarrow\infty}\alpha^{(i)}(n)$
$4$ $0.750\cdots$
$5$ $1.270\cdots$
$6$ $1.778\cdots$
$7$ $2.282\cdots$
$8$ $2.784\cdots$
We verified Conjecture 2.1   using $q$-series package in Maple. For instance, we got
$\alpha^{(4)}(65)\cdot{65^{2}}=3170.$
@@
Since the polynomial $P_{n,3}^{\,i}(q)$ is palindromic for any positive integer $i$, we can also predict the signs of the coefficients $c_{3m+1}^{(i)}(n)$ in Conjecture 2.1. This is true because of the following duality result.
Consider the polynomial $P_{n,3}^{\,i}(q)$ as defined in Conjecture 2.1. If we write
(2.1) $P_{n,3}^{\,i}(q)=X_{n,i}(q^{3})+qY_{n,i}(q^{3})+q^{2}Z_{n,i}(q^{3}),$
then it is easy to show that the following relations hold
(2.2) $X_{n,i}(q)=q^{in^{2}}X_{n,i}(1/q)$
@@
Let $n$ be a positive integer and $i\in\{2,3\}$. Consider the expansion of the polynomial $P_{n,5}^{\,i}(q)$ defined by
$P_{n,5}^{\,i}(q):=\dfrac{(q;q)_{5n}^{i}}{(q^{5};q^{5})_{n}^{i}}=\sum\limits_{m=0}^{10in^{2}}d_{m}^{(i)}(n)q^{m}.$
@@
where $\tilde{\alpha}^{(i)}(n)$ and $\tilde{\beta}^{(i)}(n)$ tend to limiting values as $n$ tends to $\infty$ according to the following table
value of $i$ $\lim\limits_{n\rightarrow\infty}\tilde{\alpha}^{(i)}(n)$ $\lim\limits_{n\rightarrow\infty}\tilde{\beta}^{(i)}(n)$
$2$ $1.133\cdots$ $1.011\cdots$
$3$ $2.132\cdots$ $2.001\cdots$
We verified Conjecture 2.2   using $q$-series package in Maple. For instance, we got
$\tilde{\beta}^{(2)}(65)\cdot{65^{2}}=4274.$
@@
Since the polynomial $P_{n,5}^{\,i}(q)$ is palindromic for any positive integer $i$, we can also predict the signs of the coefficients $d_{5m+1}^{(i)}(n)$ and $d_{5m+2}^{(i)}(n)$ in Conjecture 2.2. This is true because of the following duality result.
Consider the polynomial $P_{n,5}^{\,i}(q)$ as defined in Conjecture 2.2. If we write
(2.4) $P_{n,5}^{\,i}(q)=A_{n,i}(q^{5})+qB_{n,i}(q^{5})+q^{2}C_{n,i}(q^{5})+q^{3}D_{n,i}(q^{5})+q^{4}E_{n,i}(q^{5}),$
then it is easy to show that the following relations hold
(2.5) $A_{n,i}(q)=q^{2in^{2}}A_{n,i}(1/q),$
(2.6) $D_{n,i}(q)=q^{2in^{2}-1}C_{n,i}(1/q),$
@@
It must be noted that Wang and Krattenthaler [3, Remark $11.6.(2)$] mention that for $\delta\in\{2,3\}$, the coefficient of $q^{5m}$ in $(q;q)_{5n}^{\delta}/(q^{5};q^{5})_{n}^{\delta}$ is non-negative for all $m$, however, they do not state the limiting values of $\tilde{\alpha}^{(\delta)}(n)$ and $\tilde{\beta}^{(\delta)}(n)$.
@@ 3. Concluding Remarks
    It would be very interesting to investigate whether the asymptotic approach of Wang and Krattenthaler in [3] applies to Conjecture 2.1 and Conjecture 2.2.
@@
    The sign-pattern of coefficients of $P_{n,3}^{\,i}(q)$ is much more complicated when $i\geq 9$. For instance, we have
    $X_{11,9}(q)$ $=1-3q+8q^{3}-9q^{4}+17q^{6}-27q^{7}+46q^{9}-57q^{10}+260q^{12}$
    $+1899q^{13}+\cdots+1899q^{1076}+260q^{1077}-57q^{1079}+46q^{1080}-27q^{1082}$
    $+17q^{1083}-9q^{1085}+8q^{1086}-3q^{1088}+q^{1089},$
    where $X_{11,9}(q)$ is as defined in (2.1) with $n=11$ and $i=9$. Definitely the sign-pattern above is something worth exploring further.
  3. (3)
    Similarly, the sign-pattern of coefficients of $P_{n,5}^{\,i}(q)$ is complicated when $i\geq 4$. For instance, we have
    $\qquad\,B_{2,4}(q)$ $=-4-26q-108q^{2}-326q^{3}-748q^{4}-1536q^{5}-2644q^{6}-4352q^{7}-6200q^{8}$
    $-8418q^{9}-10544q^{10}-11369q^{11}-12772q^{12}-11134q^{13}-10072q^{14}$
    $-7762q^{15}-3696q^{16}-2348q^{17}+1420q^{18}+2281q^{19}+2924q^{20}+3422q^{21}$
    $+2300q^{22}+1983q^{23}+1044q^{24}+522q^{25}+168q^{26}-15q^{27}-48q^{28}-42q^{29}$
    $-20q^{30}-5q^{31},$
    where $B_{2,4}(q)$ is as defined in (2.4) with $n=2$ and $i=4$. Definitely the sign pattern above is again worth exploring more.
  4. (4)
    It is a challenge to find partition theoretic explanation of why $X_{n,i}(q)$ in (2.1) with $1\leq i\leq 8$ has non-negative coefficients for all $n$. Such interpretation, if any, would lead to a great insight into positivity problems in the theory of partitions and $q$-series.